OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5). −1 1 0 1/2

OpenStudy (lgbasallote):

slope = \[\frac{y_2 - y_1}{x_2 - x_1}\] therefore slope = \[\frac{-5 - (-3)}{3 - 1}\]

OpenStudy (lgbasallote):

do you get it from there?

OpenStudy (anonymous):

I just subtract from here, righht?

OpenStudy (lgbasallote):

yes you do

OpenStudy (anonymous):

I tried and im still confused, can you help me?

OpenStudy (lgbasallote):

-5 - (-3) = -5 + 3

OpenStudy (lgbasallote):

what is -5 + 3?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

except that is not a answer choice.

OpenStudy (lgbasallote):

that because -5 + 3 is not 2

OpenStudy (lgbasallote):

which is bigger..5 or 3?

OpenStudy (anonymous):

5 is bigger.

OpenStudy (lgbasallote):

and what is the sign of -5? positive or negative?

OpenStudy (anonymous):

negative

OpenStudy (lgbasallote):

therefore -5 + 3 is NEGATIVE 2 but wait that's just the numerator...look at the denominator...what's 3 - 2?

OpenStudy (anonymous):

-2/1

OpenStudy (anonymous):

i mean 1/2

OpenStudy (lgbasallote):

woops denominator is 3-1

OpenStudy (lgbasallote):

what is 3-1?

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

so \[\frac{-5 - (-3)}{3 - 1} = \frac{-2}{2}\] what does that equal??

OpenStudy (lgbasallote):

what's -2/2

OpenStudy (lgbasallote):

hey i gtg..so just divide that (use calculator if you will) whatever answer you get..that's the slope